# Taming Data and Transformers for Audio Generation

### v2

@@ 3 Method
In this section, we describe our approach to high-quality text-to-audio generation, starting with audio captioning using AutoCap in section 3.1, data collection and processing in section 3.2, and ambient audio generation with GenAu in section 3.3
@@ 3.1 Automatic Audio Captioning
Audio is an inherently ambiguous modality, as many events can produce similar sound effects—a phenomenon often leveraged in animation, where soundscapes are artificially constructed. AAC attempts to generate textual descriptions for audio clips. Previous AAC methods have generally adopted an encoder-decoder transformer design, where an audio encoder is responsible for producing a representation that is processed by the decoder to produce a caption. Recent state-of-the-art methods (Étienne Labbé et al., 2023; Kim et al., 2024b) employ a pretrained audio encoder and finetune a pre-trained language model as the decoder, relying solely on the audio content for captioning. We believe that this approach is suboptimal. By directly finetuning the pre-trained language model on the limited available dataset, these methods often suffer from overfitting and limited expressiveness and accuracy. Audio files from many sources, however, are still commonly associated with metadata that might be relevant for captioning such as raw user descriptions, or a related modality (i.e. accompanied visual information). Motivated by this observation, we propose AutoCap, an audio captioning model that employs an intermediate audio representation to connect the pretrained encoder and decoder, and uses metadata to aid with the audio captioning. Figure 1 (left) presents an overview of AutoCap.
We consider a dataset of audio signals paired with a corresponding caption $\langle\mathbf{a},\mathbf{y}\rangle$ and metadata represented as a set of token sequences $\{\mathbf{m}_{j}\}_{j=1}^{j=M}$. Inspired by state-of-the-art AAC methods (Mei et al., 2023a; Étienne Labbé et al., 2023; Kim et al., 2024b), we employ an encoder-decoder architecture. We start by computing a global feature representation of the audio:
@@
where $\mathcal{P}_{\mathrm{clap}}$ is a learnable projection layer and $\mathcal{E}_{\mathrm{clap}}$ is the audio encoder of a pretrained CLAP model (Elizalde et al., 2023). Then we compute a local feature representation of the input audio:
@@
where $\mathcal{Q}$ is a Q-Former (Li et al., 2023a) that outputs a compact sequence audio representation and $\mathcal{E}_{\mathrm{a}}$ is a pretrained HTSAT (Chen et al., 2022) audio encoder that produces a time-aligned representation. The Q-Former efficiently learns 256 latent tokens, which serve as keys in cross-attention layers with the input features, thereby condensing the audio input features into 256 tokens. Metadata sequences $\mathbf{m}_{i}$ are then embedded using the embedding layer of the pretrained decoder model to obtain corresponding embedding sequences $\mathbf{x}_{\mathrm{meta}_{i}}$. For our experiments, we use video titles and captions as the metadata. We represent the input audio and metadata as the following input sequence:
$\mathbf{x}=\mathbf{x}_{\mathrm{clap}}~{}\texttt{[boa]}~{}\mathbf{x}_{\mathrm{audio}}~{}\texttt{[eoa]}~{}\texttt{[bom]}_{1}~{}\mathbf{x}_{\mathrm{meta}_{1}}~{}\texttt{[bom]}_{1}~{}...~{}\texttt{[bom]}_{M}~{}\mathbf{x}_{\mathrm{meta}_{M}}~{}\texttt{[bom]}_{M}\text{,}$ (3)
@@
$\mathbf{\hat{y}}=\mathcal{D}_{\mathrm{t}}(\mathbf{x})\text{,}$ (4)
where $\mathcal{D}_{\mathrm{t}}$ is a pretrained BART transformer model (Lewis et al., 2020) serving as the decoder. Finally, we train our model using a standard cross-entropy loss over next token predictions:
$\mathcal{L_{\mathrm{CE}}}=-\frac{1}{T}\sum_{t=1}^{t=T}\mathrm{log}~{}p(\mathbf{y}_{t}|\mathbf{y}_{1:t-1},\mathbf{x})\text{.}$ (5)
To avoid degrading the quality of the pretrained BART and audio encoder models, we adopt a two-stage training procedure. In Stage 1, both the audio encoders and BART model are kept frozen, thus allowing the Q-Former, projection layers, and newly introduced delimiter tokens to align to the existing BART input representation. In this stage, we pretrain the model using a larger dataset of weakly-labeled audio clips. In Stage 2, we unfreeze all BART model parameters apart from the embedding layer and finetune the model on the Audiocaps dataset at a lower learning rate to make the captioning style align more closely to the target dataset. This training strategy effectively leverages the larger, weakly-labeled dataset while minimizing the knowledge drift in the pretrained BART. The use of Qformer to learn an intermediate representation is pivotal for such training strategy. Furthermore, the Qformer summarizes the audio representation into four times fewer tokens, significantly reducing the inference time.
@@ 3.2 Data Collection and Re-captioning Pipeline
Generative models in the image and video domains have shown benefits from increased quantities of data and improved quality of captions. In the audio domain, however, the major human-annotated audio-text datasets, namely AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020), provide only 51k audio clips combined. Previous methods attempted to extract additional ambient audio clips from existing video datasets using pretrained audio classifiers, but a high rejection rate of $99\%$ marked this method impractical. Instead, we found that automatic transcripts offer reliable information about the segments containing ambient sounds. In particular, we propose to select the parts of the videos that contain no automatic transcription, suggesting the absence of speech and music. Such an approach offers specific advantages over using pretrained classifiers. Automatic transcripts, readily available for most online videos, eliminate the need to download and process video and audio data before filtering. Additionally, as these transcripts provide precise time-aligned information, they facilitate the extraction of more segments, effectively reducing the rejection rate to $83\%$. Subsequently, we leverage our AutoCap model to provide textual descriptions of the extracted audio clips. Despite the effectiveness of this method in collecting ambient sounds, some clips still inadvertently contain music or speech due to transcription errors, particularly with speech in less common languages. We address this by analyzing captions and filtering out clips with keywords related to speech or music. Figure 2 summarizes our data collection pipeline.
@@
We follow this process to extract 466k audio-text pairs from Audioset (Gemmeke et al., 2017) and VGGSounds (Chen et al., 2020). Additionally, we recaption audio-only dataset such as Freesound, BBC Sound Effects, and SoundBible. To provide metadata, we employ the captioning model of Chen et al. (2024b) to extract a caption whenever a video content is available and pass an empty text otherwise. In total, we form AutoReCap, a large-scale dataset compromising of 761,113 audio-text pairs with precise captions. As an additional contribution, we introduce AutoReCap-XL, in which we scale our approach by analysing four additional large-scale video dataset (Lee et al., 2021; Xue et al., 2022; Zellers et al., 2022; Nagrani et al., 2022) with a total of 71M videos and $715.4k$ hours. In total, we collect and re-caption 57M ambient audio clips spanning $123.5k$ hours from 20.3M different videos, forming by far the largest available dataset of audio with paired captions. More details about the dataset can be found in the Appendix.
Figure 3: Scaling analysis of model size (left) and data volume with synthetic captions (right) reveal consistent improvements in FD and IS.
3.3 Scalable Text-2-Audio Generation
Latent VAE. Directly modeling waveform audio data is complex due to the high data dimensionality of audio signals. Instead, we replace the waveform with a Mel-spectrogram representation and use a VAE to further reduce its dimensionality, following prior work (Melechovsky et al., 2024; Huang et al., 2023b). Once generated, Mel-spectrograms can be decoded back to a waveform through the use of an audio vocoder (Kong et al., 2020). However, commonly-used 2D autoencoder designs (Liu et al., 2023b; c; Melechovsky et al., 2024), are not well suited to the Mel-spectrogram representation, as the separation between the Mel channels is non-linear, which is not well suited for 2D convolutions. We instead opt for a 1D-VAE design based on 1D convolutions similar to Huang et al. (2023a). We train our VAE using a combination of reconstruction, adversarial, and KL regularization losses following Esser et al. (2021).
Latent diffusion model. Following the latent diffusion paradigm, we generate audio by training a diffusion model in the latent space of the 1D-VAE. Transformer-based diffusion models currently attain state-of-the-art performance in audio generation (Huang et al., 2023a). To improve model scalability, we propose to use an efficient transformer architecture due to its success in handling long-range interactions as in video generation (Chen & Li, 2023; Menapace et al., 2024). In particular, we adopt the FIT architecture of Menapace et al. (2024) which was originally proposed to work in the pixel space and revise it for the latent space of the audio modality.
Given a 1D input $\mathbf{x}$, we first apply a projection operation to produce a sequence of input patch tokens. We then apply a sequence of FIT blocks to the input patches where each block divides patch tokens into contiguous groups of a predefined size. A set of local self-attention layers are then applied separately to each group to avoid the quadratic computational complexity of attention computation. Differently from the video domain (Menapace et al., 2024) where the high input dimensionality makes the local layers excessively expensive, we found them to be beneficial for audio generation. To further reduce the amount of computation while maintaining long-range interaction, each block considers a small set of latent tokens. First, a read operation implemented as a cross-attention layer transfers information from the patches to the latent tokens. Later, a series of global self-attention operations are applied to the latent tokens, allowing information-sharing between different groups. Finally, a write operation implemented as a cross-attention layer transfers information from the latent tokens back to the patches. Due to the reduced number of latent tokens when performing the global self-attention, computational requirements of the model are reduced with respect to a vanilla transformer design (Vaswani et al., 2017). Such a design is also particularly suited for the audio modality, which contains mostly silent or redundant parts. Unlike DiT and UNet-based methods (Ronneberger et al., 2015; Peebles & Xie, 2023) which allocate the computation resources uniformly across input tokens, the FiT architecture selectively focuses on the more informative parts.
To condition the generation on an input prompt, we use a pretrained FLAN-T5 model (Chung et al., 2022) and a CLAP (Elizalde et al., 2023) text encoder to produce the their respective embeddings $e_{\mathrm{FLAN}}$ and $e_{\mathrm{CLAP}}$, which we concatenate with the diffusion timestep $t$ to form the input conditioning signal $c$. We insert an additional cross-attention operation inside each FIT block immediately before the ‘read’ operation that makes latent tokens attend to the conditioning. Moreover, we use conditioning on dataset ID to adapt the generation style to different types of datasets.
We follow a linear noise scheduler and train the model using the epsilon prediction objective:
$\mathcal{L}=\mathbb{E}_{t,\mathbf{x},\bm{\epsilon}}~{}\big{\lVert}\mathcal{G}(\mathbf{x}_{t},c)-\bm{\epsilon}\big{\rVert}^{2}_{2}\text{,}$ (6)
where $\mathcal{G}$ is the FIT generator backbone, $\mathbf{x}_{t}$ is the input with applied noise at diffusion timestep $t$, and $\bm{\epsilon}$ is noise sampled in $N(0,1)$ with the same shape as the input.
@@ 4 Experiments
We structure the experiments section as follows: section 4.1 evaluates AutoCap by quantitatively comparing it to previous work, section 4.2 demonstrates the capabilities of GenAu quantitatively. For both, we discuss training details, baselines, metrics, results, and ablations.
4.1 Automatic Audio Captioning
Training dataset and details. We train our captioning model in two stages. During stage 1, we pretrain on a large weakly labeled dataset of 634,208 audio clips, constructed from AudioSet, Freesound, BBC Sound Effects, SoundBible, AudioCaps, and Clotho datasets. We use the ground truth captions from AudioCaps and Clotho dataset, WavCaps captions for Freesound, SoundBible, and BBC Sound Effects, and handcrafted captions through a template leveraging the provided ground truth class labels for AudioSet. As metadata, we use the title provided with each clip, and pre-compute video captions using a pretrained Panda70M model (Chen et al., 2024b) when the video modality is available or pass an empty string otherwise. We pretrain the model for 20 epochs with a learning rate of 1e-4, while keeping the audio encoder and pretrained BART frozen. In Stage 2, we fine-tune the model for 20 epochs only on AudioCaps using a learning rate of 1e-5. We randomly sample 10-second clips at 32KHz for all of our captioning experiments.
Baselines. We compare with ACT (Mei et al., 2021), V-ACT (Liu et al., 2023e), BART-tags (Gontier et al., 2021), AL-MixGEN (Kim et al., 2022), ENCLAP (Kim et al., 2024b), HTSAT-BART (Xu et al., 2023) and CNext-trans (Étienne Labbé et al., 2023). Among these baselines, ENCLAP and CNext-trans achieve the best performance. ENCLAP benefits from a stronger audio encoder and the use of a CLAP representation for additional guidance. CNext-trans trains a lightweight transformer instead of fine-tuning a pretrained language model to reduce overfitting.
Metrics and evaluation. We report results using the the established BLEU1 (Papineni et al., 2002), BLEU2 (Papineni et al., 2002), ROUGE (Lin, 2004), Meteor (Lavie & Agarwal, 2007), CIDEr (Vedantam et al., 2015), and SPIDEr (Liu et al., 2017) metrics. We evaluate our method on the AudioCaps test split using the last checkpoint of our trained model. We used only 876 clips for evaluation as some videos were deleted since the original data release. We follow the same evaluation pipeline as baselines and include their reported results. Results that were not provided in these publications are excluded from our analysis.
Results. In Tab. 1 we report the quantitative comparison. Our method outperforms previous methods on all metrics, achieving notable improvements in the CIDEr and BLUE1 scores, with values of 83.2 and 73.1, respectively. We found that incorporating metadata significantly enhances the CIDEr scores but slightly reduces the SPICE scores. This trade-off likely results from the enhanced descriptive detail brought by the metadata, which while enriching the content, introduces noise that may compromise the model’s semantic precision. In addition, AudioCaps is labeled based on audio information alone. Thus, the evaluation penalizes the description of information that can not be deduced with certainty from the audio modality only, such as the specific type of object producing a rustling sound. Compared to ENCLAP-Large (Kim et al., 2024b), and CNext-trans (Étienne Labbé et al., 2023), we find the captions produced by our method to be more descriptive and precise with a better temporal understanding. ENCLAP-Large often misses important details and exhibits lower temporal accuracy. CNext-trans, while accurate, often produces short captions that lack details. We include qualitative comparisons in the project Website. Moreover, AutoCap is four times faster than ENCALP, producing a caption for a 10-second clip in 0.28 seconds, compared to ENCALP which takes 1.12 seconds. Furthermore, we observe consistent improvements when pretraining on a large scale of weakly-labeled data during the first stage, validating the effectiveness of our training strategy in benefiting from a larger, weakly-labeled dataset.
Table 2: AutoCap ablation study on AudioCaps.
@@ 4.1 Automatic Audio Captioning
Table 3: GenAu ablation study on out-of-distribution dataset.
Model IS FD $\text{CLAP}_{MS}$
GenAU-L 18.98 20.81 0.38
GenAU-L (AC) 12.14 25.82 0.30
GenAU-S 15.76 21.29 0.36
GenAU-S w/o Recap. 11.83 25.34 0.29
Ablations. In Tab. 2, we ablate model design choices. We observe the use of the CLAP embedding to bring a $2.5$ points increase in the CIDEr score. We also validate that when not performing Stage 2 training, which involves finetuning of the BART (Lewis et al., 2020) model, performance degrades on all metrics, a finding we attribute to the necessity of adapting BART’s decoder to the sentence structure typical of AudioCaps. A more severe degradation in performance is observed if Stage 1 is not performed, with the misaligned representation between the encoder and the decoder causing catastrophic forgetting in the language model. Finally, if BART word embeddings are finetuned in Stage 2 instead of being kept frozen, we observe a slight performance degradation.
4.2 Text-2-Audio Generation
Training dataset and details. We train on similar data settings to baselines. We use our best-performing captioning model to re-caption the WavCaps dataset. In addition, we obtain 339,387 videos from AudioSet and 126,905 videos from VGGSounds, totaling 761,113 clips. For those obtained from sound-only platforms, we input an empty string as the video caption. For full details of the data sources of our training dataset, please refer to the Appendix. We additionally use Clotho and AudioCaps training datasets with their ground truth caption. To stay consistent with baselines, we train at 16kHz resolution. We use a patch size of 1 and a group size of 32. We use LAMB optimizer (You et al., 2020) with a LR of 5e-3. We train for 220k steps and choose the checkpoint with the highest IS, at steps 210k and 207k for the large and small models. We also disable EMA as found it to make the metrics unstable.
Baselines. We compare with TANGO 1 & 2, (Ghosal et al., 2023), AudioLDM 1 & 2 (Liu et al., 2023b; c), and Make-An-Audio 1 & 2 (Huang et al., 2023b; a). Both AudioLDM and Make-an-Audio train a UNet-based latent diffusion model (Rombach et al., 2022) on Mel-Spectrogram representation of the audio, by regarding the Mel-Spectrogram as a single channel image, and use a pretrained CLAP encoder to condition the generation on an input prompt. TANGO proposed to use FLAN-T5 (Chung et al., 2022) as the text encoder and reported significant improvements. AudioLDM-2 and Make-an-Audio-2 proposed to use a dual encoder strategy of a T5 (Raffel et al., 2022) and CLAP encoder. AudioLDM-2 focused on extending the generation and conditioning to various domains. Specifically, they use the language of audio (LOA) to condition the generation on images, audio, or transcripts and train their model for music and speech generation. Make-an-Audio-2 proposes to use a 1D VAE representation and employ a feed-forward Transformer-based model to replace the UNet. Recently, Tango-2 proposed to use instruction fine-tuning on a synthetic dataset to enhance the temporal understanding. In our experiments, we focus on text-conditioned natural audio generation and generate 10s clips at a resolution of 16Khz.

### v3

@@ 3 Method
In this section, we describe our approach to high-quality text-to-audio generation, starting with audio captioning using AutoCap in section 3.1, data collection in section 3.2, and ambient audio generation with GenAu in section 3.3
@@ 3.1 Automatic Audio Captioning
Recent state-of-the-art methods (Labb et al., 2024; Kim et al., 2024b) generally employ an encoder-decoder transformer design where a pretrained audio encoder passes the audio representation to a pre-trained language model serving as the decoder. This language model (e.g. BART) is typically finetuned to adapt to the audio representation. However, due to the distribution mismatch between the pretraining data of the LLM and the audio embeddings produced by the encoder, the decoder suffers from catastrophic forgetting. Furthermore, audio is an inherently ambiguous modality, as many events can produce similar sound effects—a phenomenon often leveraged in animation, where soundscapes are artificially constructed. Audio clips from many sources, however, are still commonly associated with metadata that might be relevant for captioning such as raw user descriptions, or related modalities (i.e. accompanied visual information). Motivated by these observations, we propose AutoCap, an audio captioning model that employs an intermediate audio representation to connect the pretrained encoder and decoder and uses metadata to aid with the captioning. Figure 1 (left) presents an overview of AutoCap.
We consider a dataset of audio-caption pairs $\langle\mathbf{a},\mathbf{y}\rangle$ and corresponding metadata represented as a set of token sequences $\{\mathbf{m}_{j}\}_{j=1}^{j=M}$. Inspired by state-of-the-art AAC methods (Mei et al., 2024a; Labb et al., 2024; Kim et al., 2024b), we employ an encoder-decoder architecture. We first compute a global feature representation of the audio:
@@
where $\mathcal{P}_{\mathrm{clap}}$ is a learnable projection layer and $\mathcal{E}_{\mathrm{clap}}$ is the audio encoder of a pretrained CLAP model (Wu et al., 2023a). We also compute local features as:
@@
where $\mathcal{Q}$ is a Q-Former (Li et al., 2023) and $\mathcal{E}_{\mathrm{a}}$ is a pretrained HTSAT (Chen et al., 2022) audio encoder that produces a time-aligned representation (1024 tokens) following (Mei et al., 2024a). The Q-Former efficiently learns 256 latent tokens, which serve as keys in cross-attention layers with the input features, thereby outputting 256 tokens. Metadata sequences $\mathbf{m}_{i}$ are then embedded using the embedding layer of a pretrained BART to obtain embedding sequences $\mathbf{x}_{\mathrm{meta}_{i}}$. For our experiments, we use video titles and captions as the metadata. We represent the input audio and metadata as the following input sequence:
$\mathbf{x}&=\mathbf{x}_{\mathrm{clap}}~{}\texttt{\small[boa]}~{}\mathbf{x}_{\mathrm{audio}}~{}\texttt{\small[eoa]}~{}\texttt{\small[bom]}_{1}~{}\mathbf{x}_{\mathrm{meta}_{1}}\\ &\quad~{}\texttt{\small[bom]}_{1}~{}...~{}\texttt{\small[bom]}_{M}~{}\mathbf{x}_{\mathrm{meta}_{M}}~{}\texttt{\small[bom]}_{M}\end{aligned}\text{,}$ (3)
@@
Training. We train our model using a standard cross-entropy loss over next token predictions. To avoid degrading the quality of the pretrained BART and audio encoder models, we adopt a two-stage training procedure. In Stage 1, both the audio encoders and BART model are kept frozen, thus allowing the Q-Former, projection layers, and newly introduced delimiter tokens to align to the pretraining BART representation. In this stage, we pretrain the model using a larger dataset of weakly-labeled audio clips. In Stage 2, we unfreeze all BART parameters apart from the embedding layer and finetune the model on the Audiocaps dataset at a lower learning rate to align the captioning style more with human style. This training strategy leverages the larger, weakly-labeled dataset while minimizing the knowledge drift in the pretrained BART. The use of Q-Former to learn an intermediate representation is pivotal for such a strategy.
Figure 3: Scaling analysis of model size (left) and data with synthetic captions (right) reveal consistent improvements in FD and IS.
@@ 3.2 Data Collection and Re-captioning Pipeline
Generative models in the image and video domains have shown benefits from increased quantities of data and improved quality of captions. In the audio domain, however, the major human-annotated audio-text datasets, namely AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020), provide only 51k audio clips combined. Previous methods attempted to extract additional ambient audio clips from existing video datasets using pretrained audio classifiers, but a high filtering rate marked this method impractical. Instead, we found that automatic transcripts offer reliable information about the segments containing ambient sounds. In particular, we propose to select the parts of the videos that contain no automatic transcription, suggesting the absence of speech and music. Such an approach offers specific advantages over using pretrained classifiers. Automatic transcripts, readily available for most online videos, eliminate the need to download and process video and audio data before filtering. Additionally, as these transcripts provide precise time-aligned information, they facilitate the extraction of more segments per video. Subsequently, we leverage our AutoCap model to provide textual descriptions of the extracted audio clips. Despite the effectiveness of this method in collecting ambient sounds, some clips still inadvertently contain music or speech due to transcription errors, particularly with speech in less common languages. We address this by analyzing captions and filtering out clips with keywords related to speech or music. Finally, we filter all audio-text pairs with CLAP similarity less than 0.1.
@@
We follow this process to extract 466k audio-text pairs from Audioset (Gemmeke et al., 2017) and VGGSounds (Chen et al., 2020). Additionally, we recaption audio-only datasets such as Freesound, BBC Sound Effects, and SoundBible. To provide metadata, we employ the captioning model of Chen et al. (2024b) to extract a caption whenever video content is available and pass an empty text otherwise. In total, we form AutoReCap, a large-scale dataset compromising of 761,113 audio-text pairs with precise captions. As an additional contribution, we introduce AutoReCap-XL, in which we scale our approach by analyzing four additional large-scale video dataset (Lee et al., 2021; Xue et al., 2022; Zellers et al., 2022; Nagrani et al., 2022) with a total of 71M videos and $715.4k$ hours. After filtering, we collect and re-caption 47M ambient audio clips spanning $123.5k$ hours from 20.3M different videos, forming by far the largest available dataset of audio with paired captions. Figure 2 summarizes our data collection pipeline, and Sec. A in Appendix present more details about the dataset collection and processing.
@@
3.3 Scalable Text-2-Audio Generation
Latent VAE. Directly modeling waveforms is complex due to the high data dimensionality of audio signals. Instead, we replace the waveform with a Mel-spectrogram representation and use a VAE to further reduce its dimensionality, following prior work (Melechovsky et al., 2023; Huang et al., 2023b). Once generated, Mel-spectrograms can be decoded back to a waveform through a vocoder (Lee et al., 2022). However, commonly-used 2D autoencoder designs (Liu et al., 2023a, b; Melechovsky et al., 2023), are not well suited to the Mel-spectrograms, as the separation between the Mel channels is non-linear, which is not well suited for 2D convolutions. We instead opt for a 1D-VAE design based on 1D convolutions similar to Huang et al. (2023a). We train the VAE following Esser et al. (2021).
Latent diffusion model. Following the latent diffusion paradigm, we generate audio by training a diffusion model in the latent space of the 1D-VAE. Transformer-based diffusion models currently attain state-of-the-art performance in audio generation (Huang et al., 2023a). However, both UNet and transformer-based baselines exhibited limited performance gains with increasing model size (Liu et al., 2023b; Hai et al., 2024). We observe that ambient audio often contains extensive silent and redundant segments, which may explain the poor scalability of UNet and DiT-based methods, as they distribute computation uniformly across the input. Therefore, we propose to use a more dynamic transformer architecture as a diffusion backbone (Chen & Li, 2023; Menapace et al., 2024). In particular, we adopt the FIT architecture of Menapace et al. (2024) which was originally proposed to work in the pixel space, and revise it for the latent space of the audio modality.
Given a 1D input $\mathbf{x}$, we first apply a projection operation to produce a sequence of input patch tokens. We then apply a sequence of FIT blocks to the input patches where each block divides patch tokens into contiguous groups of a predefined size. A set of local self-attention layers are then applied separately to each group to avoid the quadratic computational complexity of attention computation. Differently from the video domain (Menapace et al., 2024) where the high input dimensionality makes the local layers excessively expensive, we found them to be beneficial for audio generation. To further reduce the amount of computation while maintaining long-range interaction, each block considers a small set of latent tokens. First, a read operation implemented as a cross-attention layer transfers information from the patches to the latent tokens. Later, a series of global self-attention operations are applied to the latent tokens, allowing information-sharing between different groups. Finally, a write operation implemented as a cross-attention layer transfers information from the latent tokens back to the patches. Due to the reduced number of latent tokens when performing the global self-attention, computational requirements of the model are reduced with respect to a vanilla transformer design (Vaswani et al., 2017). Such a design is particularly suited for the audio modality, which contains mostly silent or redundant parts. Unlike DiT and UNet-based methods (Ronneberger et al., 2015; Peebles & Xie, 2023) which allocate the computation resources uniformly across input tokens, the FTT architecture selectively focuses on the more informative parts, dedicating more compute for these parts as the model size scales.
To condition the generation on an input prompt, we use a pretrained FLAN-T5 model (Chung et al., 2024) and a CLAP (Wu et al., 2023a) text encoder to produce the their respective embeddings $e_{\mathrm{FLAN}}$ and $e_{\mathrm{CLAP}}$ following prior work (Liu et al., 2023b), which we concatenate with the diffusion timestep $t$ to form the input conditioning signal $c$. We insert an additional cross-attention operation inside each FIT block immediately before the ‘read’ operation that makes latent tokens attend to the conditioning. Moreover, we use conditioning on dataset ID to adapt the generation style to different datasets. We train the model using the epsilon prediction objective and follow a linear noise scheduler.
@@ 4 Experiments
Table 2: AutoCap ablation study on AudioCaps
@@
In section 4.1, we evaluate AutoCap quantitatively. We then demonstrate the capabilities of GenAu in section 4.2 and discuss scaling trends with respect to data and model size. We also provide qualitative comparisons on the Website.
4.1 Automatic Audio Captioning
Training dataset and details. We train AutoCap in two stages. During stage 1, we pretrain on a large weakly labeled dataset of 634,208 audio clips, constructed from AudioSet, Freesound, BBC Sound Effects, SoundBible, AudioCaps, and Clotho. We use ground truth captions from AudioCaps and Clotho datasets, WavCaps captions for Freesound, SoundBible, and BBC Sound Effects, and handcrafted captions through a template leveraging the ground truth class labels for AudioSet. As metadata, we use the title provided with each clip, and pre-compute video captions using a pretrained Panda70M model (Chen et al., 2024b) or pass an empty string when the video modality is unavailable. We pretrain the model for 20 epochs with a learning rate of 1e-4, while keeping the audio encoder and pretrained BART frozen. In Stage 2, we fine-tune the model for 20 epochs on AudioCaps using a learning rate of 1e-5. We use 10-second clips at 32KHz for all experiments.
Baselines. We compare with V-ACT (Liu et al., 2023c), BART-tags (Gontier et al., 2021), AL-MixGEN (Kim et al., 2022), ENCLAP (Kim et al., 2024b), HTSAT-BART (Xu et al., 2024b), CNext-trans (Labb et al., 2024) and GAMA (Ghosh et al., 2024b). Among these, ENCLAP and CNext-trans have the best performance. ENCLAP benefits from a stronger audio encoder and a CLAP representation. CNext-trans trains a lightweight transformer instead of fine-tuning a pretrained language model to reduce overfitting.
Metrics and evaluation. We report results using the established BLEU1 and BLEU4 (Papineni et al., 2002), ROUGE (Lin, 2004), Meteor (Lavie & Agarwal, 2007), CIDEr (Vedantam et al., 2015), and SPIDEr (Liu et al., 2017) metrics. We evaluate our method on the AudioCaps test split using the last checkpoint of our trained model. We follow the same evaluation pipeline as baselines and include their reported results, except for GAMA which we evaluate using their released checkpoint. Metrics unavailable in these publications are excluded from our analysis.